# Can Helicity Modulus Be Defined For Boundary Conditions With Finite Twist?

### v3

@@ V Conclusion
To summarize, in this paper we have studied the effects of finite (non-infinitesimal) twists in the boundary conditions on a two-dimensional classical ferromagnetic XY model. The system’s response to the twist has been studied by a direct evaluation of the free energy by means of large-scale Monte Carlo simulations. We have found that in the quasi long-range ordered phase below the BKT transition, the free energy cost of a non-infinitesimal twist deviates from the expected quadratic dependence on the twist angle. In the case of a $\pi$ twist (anti-periodic boundary conditions), the mixing of states of opposite chiralities causes an excess entropy contribution of $\ln(2)$ that lowers the free energy cost of the twist. Thus, if the helicity modulus is calculated from the free energy of a $\pi$-twisted system (using eq. (4)), its value is reduced compared to the usual helicity modulus obtained from an infinitesimal twist. We note in passing that our discussion has been based on the phase representation of the XY model, but the same results are expected in other representations [40, 41, 42, 43] provided they correctly reflect the chiral symmetry of the boundary conditions.
@@
Chirality mixing is important for computer simulations for at least two reasons. First, efficient simulations often involve nonlocal algorithms (such as the Wolff cluster algorithm) that can freely flip the bulk chirality of the system. Moreover, system sizes in computer simulations are sometimes not very large so that even algorithms with local updates (such as the Metropolis algorithm) may allow the system to fluctuate between different chiralities. Free energies of $\pi$-twisted systems computed in such simulations contain the $-T\ln(2)$ entropic free energy contribution due to chirality mixing which needs to corrected to compute the usual (infinitesimal-twist) helicity modulus $\rho_{s0}$ from simulations with a $\pi$ twist. When determining the BKT transition temperature from the helicity modulus data, one can, alternatively, modify the Kosterlitz-Nelson relation as discussed in Sec. IV.1.
The partition function of a quantum system in $d$ dimensions can often be mapped onto that of a classical system in $d+1$ dimensions. Our findings imply that a macroscopic one-dimensional quantum rotor model with antiperiodic boundary conditions has a doubly degenerate ground state because the tunneling time between states of opposite chirality diverges with system size.
@@
Our work also relates to some of the questions raised by Brown and Ciftan [15]. They studied the effects of twisted-periodic and anti-periodic boundary conditions on a three-dimensional classical Heisenberg model and discussed the notion of mixing states of different chiralities in response to twisted-periodic boundary conditions. However, they analyzed the internal energy cost of the twist rather than the free energy cost. The authors report deviations from a quadratic twist angle dependence of the internal energy cost somewhat similar to what we find for $\Delta F$, but the magnitude of the deviation in their case is much larger than $T\ln(2)$. Moreover, the internal energy cost (as opposed to $\Delta F$) is not expected to contain the entropy due to the mixing of chiralities. This suggests that the deviations from a quadratic twist angle dependence of the internal energy cost in Ref. [15] have a different origin. Specifically, Heisenberg spins can reduce the energy cost of twisted-periodic boundary conditions by tilting out of the plane in which the twist is applied. In contrast, they cannot avoid antiperiodic boundary conditions, in agreement with the fact that the energy cost of anti-periodic boundary conditions in Ref. [15] is much larger that of even the largest twist angles. A quantitative analysis of the effects of twisted-periodic boundary conditions in the Heisenberg case remains a task for the future.

### v4

@@ V Conclusion
To summarize, in this paper we have studied the effects of finite (non-infinitesimal) twists in the boundary conditions on a two-dimensional classical ferromagnetic XY model. The system’s response to the twist has been studied by a direct evaluation of the free energy by means of large-scale Monte Carlo simulations. We have found that in the quasi long-range ordered phase below the BKT transition, the free energy cost of a non-infinitesimal twist deviates from the expected quadratic dependence on the twist angle. In the case of a $\pi$ twist (anti-periodic boundary conditions), the mixing of states of opposite chiralities causes an excess entropy contribution of $\ln(2)$ that lowers the free energy cost of the twist. Thus, if the helicity modulus is calculated from the free energy of a $\pi$-twisted system (using eq. (4)), its value is reduced compared to the usual helicity modulus obtained from an infinitesimal twist. We note that our results have been obtained by comparing periodic and twisted-periodic boundary conditions. However, all our arguments are expected to hold as well for systems in which the boundary spins are held at fixed angles. We also note that our discussion has been based on the phase representation of the XY model, but the same results are expected in other representations [42, 43, 44, 45] provided they correctly reflect the chiral symmetry of the boundary conditions.
@@
Chirality mixing is important for computer simulations for at least two reasons. First, efficient simulations often involve nonlocal algorithms (such as the Wolff cluster algorithm) that can freely flip the bulk chirality of the system. Moreover, system sizes in computer simulations are sometimes not very large so that even algorithms with local updates (such as the Metropolis algorithm) may allow the system to fluctuate between different chiralities. Free energies of $\pi$-twisted systems computed in such simulations contain the $-T\ln(2)$ entropic free energy contribution due to chirality mixing. To compute the usual (infinitesimal-twist) helicity modulus $\rho_{s0}$ from simulations with a $\pi$ twist, this entropic contribution needs to be removed. When determining the BKT transition temperature from the helicity modulus data, one can, alternatively, modify the Kosterlitz-Nelson relation as discussed in Sec. IV.1.
The effects of twisted boundary conditions on an XY model in thin-film geometry were also studied in Ref. [13] using a mean-field theory. For a twist of $\Theta=\pi$, this work finds an additional singularity of the free energy below the bulk critical temperature. It is induced by the boundary conditions and leads to kink in the Casimir force. In our calculations, we do not observe such a singularity. We believe that this may stem from the fact that the mechanism proposed in Ref. [13], viz., a competition between a rotational state and a planar state to fulfill the boundary conditions, requires soft spin variables and does not hold for the hard spins (of fixed unit length) of our XY model. Alternatively, the lack of additional singularity might be because states with both directions of rotation are simultaneously realized in our simulations at all twist angles, whereas in Ref. [13], only one direction is considered in the calculations.
The partition function of a quantum system in $d$ dimensions can often be mapped onto that of a classical system in $d+1$ dimensions. Our results thus apply to a one-dimensional (particle-hole symmetric) quantum rotor model with twisted boundary conditions. Specifically, our findings imply that in a finite-size rotor model the states with opposite chirality hybridize. However in the thermodynamic limit, the ground state is doubly degenerate because the hybridization goes to zero and the tunneling time between states of opposite chirality diverges with system size. Analogous results are expected in a field-theoretic approach based on the sine-Gordon model [46, 47].
@@
Our work also relates to some of the questions raised by Brown and Ciftan [17]. They studied the effects of twisted-periodic and anti-periodic boundary conditions on a three-dimensional classical Heisenberg model and discussed the notion of mixing states of different chiralities in response to twisted-periodic boundary conditions. However, they analyzed the internal energy cost of the twist rather than the free energy cost. The authors report deviations from a quadratic twist angle dependence of the internal energy cost somewhat similar to what we find for $\Delta F$, but the magnitude of the deviation in their case is much larger than $T\ln(2)$. Moreover, the internal energy cost (as opposed to $\Delta F$) is not expected to contain the entropy due to the mixing of chiralities. This suggests that the deviations from a quadratic twist angle dependence of the internal energy cost in Ref. [17] have a different origin. Specifically, Heisenberg spins can reduce the energy cost of twisted-periodic boundary conditions by tilting out of the plane in which the twist is applied. In contrast, they cannot avoid antiperiodic boundary conditions, in agreement with the fact that the energy cost of anti-periodic boundary conditions in Ref. [17] is much larger than that of even the largest twist angles. A quantitative analysis of the effects of twisted-periodic boundary conditions in the Heisenberg case remains a task for the future.
@@ Acknowledgements.
Appendix A Helicity modulus for an infinitesimal twist
In this Appendix, we derive the expression (5) for the second derivative of the free energy on the twist angle. This expression allows the evaluation of helicity modulus for an infinitesimal twist in terms of correlation functions of the untwisted system. The derivation follows Refs. [25, 50].
Consider an XY magnet in $d$ dimensions, of linear size $L$. Applying twisted boundary conditions (BC) along $x$ axis implies that the spins at $x=0$ make an angle $\Theta$ w.r.t. the spins at $x=L$. The change in free energy due to the twist can be parameterized as
$\Delta F=\frac{1}{2}\rho_{s}V\left(\frac{\Theta}{L}\right)^{2},$ (17)
where $V=L^{d}$. In the limit of $\Theta\rightarrow 0$,
$\rho_{s}=\frac{L^{2}}{V}\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}.$ (18)
The second derivative of the free energy can be evaluated by treating $\Theta$ as a parameter of the partition function.
To this end, we start from the Hamiltonian
$H=-\sum_{<ij>}J\cos(\phi_{i}-\phi_{j})$ (19)
with boundary conditions $\phi_{i}=0$ at $x_{i}=0$ and $\phi_{i}=\Theta$ at $x_{i}=L$. We now perform a variable transformation $\psi_{i}=\phi_{i}-\Theta x_{i}/L$. Note that $\psi_{i}$ has untwisted BC, $\psi_{i}=\phi_{i}=0$ at $x_{i}=0$ and $x_{i}=L$. The dependence on the twist angle $\Theta$ has been moved to the Hamiltonian
$H=-\sum_{<ij>}J\cos(\psi_{i}+\Theta\frac{x_{i}}{L}-\psi_{j}-\Theta\frac{x_{j}}{L})~{},$ (20)
$H=-\sum_{<ij>}J\cos(\psi_{i}-\psi_{j}-\frac{\Theta}{L}(x_{i}-x_{j}))~{}.$ (21)
The evaluation of $(\partial^{2}F/\partial\Theta^{2})$ is now straight forward, starting from
$F=-k_{B}T\ln(Z({\Theta}))=-k_{B}T\ln\left(\textrm{Tr}e^{-\beta H}\right)~{}.$ (22)
First,
$\frac{\partial F}{\partial\Theta}=$ $-k_{B}T\frac{1}{Z}\left(\frac{\partial Z}{\partial\Theta}\right)$
$\frac{\partial F}{\partial\Theta}=$ $\frac{1}{Z}\textrm{Tr}\left(\frac{\partial H}{\partial\Theta}e^{-\beta H}\right)=\left\langle\frac{\partial H}{\partial\Theta}\right\rangle.$
Then,
$\frac{\partial^{2}F}{\partial\Theta^{2}}$ $=\frac{\partial}{\partial\Theta}\left(\frac{1}{Z}\textrm{Tr}\left(\frac{\partial H}{\partial\Theta}e^{-\beta H}\right)\right)$
$\frac{\partial^{2}F}{\partial\Theta^{2}}$ $=\beta\left\langle\frac{\partial H}{\partial\Theta}\right\rangle^{2}+\left\langle\frac{\partial^{2}H}{\partial\Theta^{2}}\right\rangle-\beta\left\langle\left(\frac{\partial H}{\partial\Theta}\right)^{2}\right\rangle.$
Each derivative in the above equation can be evaluated from Eq. (21).
$\left(\frac{\partial H}{\partial\Theta}\right)_{\Theta=0}=$ $\frac{1}{L}\sum_{<ij>}J\sin(\psi_{i}-\psi_{j})(x_{i}-x_{j})$ (23)
$\left(\frac{\partial^{2}H}{\partial\Theta^{2}}\right)_{\Theta=0}=$ $\frac{1}{L^{2}}\sum_{<ij>}J\cos(\psi_{i}-\psi_{j})(x_{i}-x_{j})^{2}.$ (24)
Collecting all the terms,
$\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}=\beta\left\langle\frac{\partial H}{\partial\Theta}\right\rangle^{2}_{\Theta=0}+\left\langle\frac{\partial^{2}H}{\partial\Theta^{2}}\right\rangle_{\Theta=0}-\beta\left\langle\left(\frac{\partial H}{\partial\Theta}\right)^{2}\right\rangle_{\Theta=0}.$ (25)
The first term vanishes due to symmetry, but
$\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}=$ $\frac{1}{L^{2}}\left\langle\sum_{<ij>}J\cos(\psi_{i}-\psi_{j})(x_{i}-x_{j})^{2}\right\rangle$
-- $\frac{\beta}{L^{2}}\left\langle\sum_{<ij>}\left\{J\sin(\psi_{i}-\psi_{j})(x_{i}-x_{j})\right\}^{2}\right\rangle$
where the thermodynamic averages delimited-⟨$\langle$..$\rangle$ … ⟩ are defined for $\Theta=0$. In vector notation, $\cos(\psi_{i}-\psi_{j})=\vec{S}_{i}\cdot\vec{S}_{j},~{}\sin(\psi_{i}-\psi_{j})=-\hat{k}\cdot(\vec{S}_{i}\times\vec{S}_{j})$. Thus,
$~{}~{}~{}~{}\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}$ $=\frac{1}{L^{2}}\sum_{<ij>}J\langle\vec{S}_{i}\cdot\vec{S}_{j}\rangle(x_{i}-x_{j})^{2}$
$-\frac{\beta}{L^{2}}\left\langle\left\{\sum_{<ij>}J(\hat{k}\cdot(\vec{S}_{i}\times\vec{S}_{j})(x_{i}-x_{j}))\right\}^{2}\right\rangle.$
The $\Theta\rightarrow 0$ limit thus allows a simple evaluation of $({\partial^{2}F}/{\partial\Theta^{2}})$ and of the helicity modulus $\rho_{s}$. This completes the derivation of Eq. (5).